An experiment has 6 different outcomes. The probabilities of the first five outcomes are 3/20, 7/20,...

  1. An experiment has 6 different outcomes. The probabilities of the first five outcomes are 3/20, 7/20, 1/10, 1/25 and 4/25. What is the probability of the last outcome?
Answer #1

For probability, the sum of all the outcomes = 1

Let the last outcome = x

Therefore (3/20) + (7/20) + (1/10) + (1/25) + (4/25) + x = 1

(3/20) + (7/20) + (1/10) + (1/25) + (4/25) = (15 + 35 + 10 + 4 + 16) / 100 = 80 / 100 = 4 / 5

Therefore the 6th outcome = 1 - 4/5 = 1/5